OpenStudy (anonymous):

IF PA= 10 and BE = 21 find BP

OpenStudy (anonymous):

the question is not clear. attach an image to go with it.

OpenStudy (anonymous):

If PA is 10 and BE is 21 then what is BP

OpenStudy (anonymous):

if BE = 21, then the radius is 21/2 say the center of the circle is Z then AZ^2+PA^2 = PZ^2 find PC then PB = PZ+ZB ZB is the radius

OpenStudy (anonymous):

this is possible because the tangent to a circle is perpendicular to the line joining the tangent and the center of the circle. so we use the pythagoras theorem